De-broglie wavelength of electron is related to applied voltage is 
 

Grade:11

2 Answers

Khimraj
5 years ago
λde Broglie=
h/p = h/me.ve =
since p2/2m = E = eV
so p = \sqrt{2meV}
so \lambda = h/\sqrt{2meV}
…............................................................
Khimraj
5 years ago
\lambda = h/p
p2/2m = E = eV
so p = \sqrt{2meV}
so \lambda  = h/\sqrt{2meV}
Hope it clears........................................................................................................